# The Memorization Problem: Can We Trust LLMs’ Economic Forecasts?

### v1

@@ 2 Methodology
Our testing spans four categories of economic variables. First, we examine stock market indices through questions about exact numerical levels, directional changes, percentage movements, and relative performance. For example, we ask for the S&P 500 closing value on specific dates or whether the NASDAQ increased or decreased on particular days. Second, we test macroeconomic indicators by querying precise values (e.g., unemployment rates) and directional trends. Third, we assess individual securities information, including specific stock price levels and directional movements. Fourth, we test LLMs’ ability to identify news headlines’ dates by presenting the model with sets of The Wall Street Journal front page headlines (without dates) and asking it to identify when these headlines appeared, as well as to predict the corresponding S&P 500 level on the following trading day.
@@ 2.1 Period Selection and Supplementary Analysis
Our experimental design strategically spans three temporal zones relative to each model’s training cutoff date. First, we include periods before the cutoff where we expect high recall accuracy if memorization occurs. Second, we test periods within 10 years of the cutoff to examine potential recency effects in memorization patterns. Finally, we include post-cutoff periods as a control condition where memorization is impossible.
@@ 2.2 Prompt
The context information represents the information set $x_{t}$ for our experimental conditions, which may be empty (testing pure recall), contain historical data points, include relevant news headlines, or provide general knowledge about the period.
For example, in a prompt testing recall with historical context, we might provide: ”Context: The S&P 500 closed at 2,834.40 on March 14, 2019, and at 2,808.48 on March 13, 2019. What was the S&P 500 closing value on March 15, 2019?” This standardized approach allows us to systematically vary the information provided while controlling for confounding factors in question phrasing or response expectations.
2.3 LLM
Given its wide usage in research, we run our analysis using ChatGPT 4o with the specific version of “gpt-4o-2024-08-06,” whose training data ends in October 2023.
@@ 3 Data
Most of the tests use data from January 1990 to September 2023. To test the LLM’s memorization, we use three categories of datasets: (1) stock index and individual stock prices, (2) macro indicators, and (3) textual data.
We ask the LLM to give us the closing value of the stock indices and a sample of individual stocks. We use the daily closing values of the S&P 500, the Dow Jones Industrial Average, and the Nasdaq Composite from Yahoo Finance to evaluate the LLM’s answers. We use Center for Research in Security Prices (CRSP) data to obtain daily stock market data for individual stock closing prices. The sample of individual stocks includes the Magnificent 7 (AAPL, AMZN, GOOGL, META, MSFT, NVDA, TSLA) and ten randomly selected from each of the small, mid, and large-cap categories. We resample the random stocks chosen yearly to account for changes in size over time.
We also ask the LLM to give us estimates of various macroeconomic indicators. The indicators we test are US GDP growth, inflation, unemployment rate, 10-year Treasury Yield, VIX, housing starts, and change in nonfarm payrolls. We obtain the actual unemployment rate and 10-year Treasury Yield values from Federal Reserve Economic Data (FRED). We obtain the VIX levels from Yahoo Finance. For GDP growth, inflation, housing starts, and change in nonfarm payrolls, we use the Philadelphia Federal Reserve Real-Time Data Set to get the first vintage and ask the LLM to give us the earliest estimate of these indicators.
[…]
@@ 4 Results
In this section, we present a comprehensive evaluation of GPT-4o’s memorization of economic and financial data, spanning headlines, macroeconomic indicators, market indices, individual stocks, portfolios by market cap, and attempts to mitigate memorization through knowledge cutoffs and masking. Across these domains, we assess the model’s ability to recall precise values, identify contextual details, and adhere to constraints. We use data from 1989 to 2025, with pre-cutoff and post-cutoff periods (October 2023), to distinguish memorization from inference. Our findings reveal the extent and selectivity of memorization, highlighting its implications for using LLMs in financial forecasting and the challenges of isolating genuine predictive ability.
Each subsection examines a specific data type or mitigation strategy, building a cohesive picture of how memorization manifests and persists. From the near-perfect recall of headline dates and macroeconomic rates to the varying accuracy for stock prices by firm prominence and the limitations of cutoff instructions and masking, our results underscore a consistent pattern: GPT-4o’s performance on pre-cutoff data often reflects training data exposure rather than analytical insight. These analyses collectively inform our understanding of the risks in historical financial research with LLMs and the need for rigorous evaluation beyond training boundaries.
@@ 4.1 Macroeconomic Data Recall
To assess GPT-4o’s memorization of macroeconomic indicators, we tested its ability to recall monthly values (quarterly for GDP) across various variables, using data from January 1990 to September 2023, all within the model’s training cutoff of October 2023. The indicators were divided into two groups: rates (GDP Growth, Inflation, Unemployment Rate, and 10-year Treasury Yield) and levels (Housing Starts, VIX, and Nonfarm Payrolls). For rates, we requested percentage values, evaluating accuracy through Mean Error, Mean Absolute Error, Directional Accuracy (correctly identifying whether the rate was above a threshold value), and Directional Accuracy Change (correct direction of change from the previous period). For levels, we requested raw values, with errors calculated relative to actual levels through Mean Percent Error, Mean Absolute Percent Error, Directional Accuracy, and Directional Accuracy Change. We also examined levels over a recent 10-year period to explore potential recency effects. Performance was measured against actual values from Federal Reserve Economic Data (FRED), Yahoo Finance, and the Philly Fed Real-Time data, with results reported in Table 1 and Figure 3.
@@
The results reveal an evident ability to recall macroeconomic data. For rates, the model demonstrates near-perfect recall, with Mean Absolute Errors ranging from 0.03% (Unemployment Rate) to 0.15% (GDP Growth) and Directional Accuracy exceeding 96% across all indicators, reaching 98% for 10-year Treasury Yield and 99% for Unemployment Rate. This result suggests that GPT-4o has memorized these percentage-based indicators with high fidelity.
@@
For levels, the recall remains high, with Directional Accuracies between 92% and 100% for all indicators during the whole pre-training sample. Moreover, when focusing on the most recent 10-year period in the pre-training sample, performance improves dramatically–Mean Absolute Percent Errors fall to 1.06% for Housing Starts, 0.34% for VIX, and below 0.00% for Nonfarm Payrolls, with Directional Accuracy rising to 95%–100%. This recency effect indicates stronger memorization for more recent data, possibly due to denser representation in the training corpus.
@@
The high recall accuracy for rates and recent levels underscores the memorization problem when evaluating LLMs’ forecasting capabilities. The model’s ability to reproduce precise macroeconomic values, especially for percentage-based indicators and recent periods, suggests that apparent forecasting success for pre-cutoff data may stem from retrieving memorized figures rather than genuine economic analysis. The weaker performance for levels over the full period, particularly for volatile indicators like Nonfarm Payrolls, hints at selective memorization, where certain data types or time frames are less reliably retained. These findings reinforce concerns about using LLMs to analyze historical economic data, as their outputs may reflect training data exposure rather than predictive insight.
@@
This table reports a set of evaluation metrics for various macroeconomic indicators grouped into three panels: Rates, Levels, and Levels, Recent Period: Past 10 years. We ask the LLM to recall monthly values (quarterly for GDP, specific end of month date for 10-Year Treasury Yield and VIX) for each indicator. The indicators in the Rates panel include GDP Growth, Inflation, Unemployment Rate, and the 10-Year Treasury Yield. For these indicators, we ask the LLM to give us a percentage. The Levels panel includes Housing Starts, VIX, and Nonfarm Payrolls, evaluated over the full sample period. The Levels, Recent Pre-cutoff Period: Past 10 years panel evaluates these same indicators over a more recent, shorter period. Mean Error (ME), Mean Absolute Error (MAE), Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), Directional Accuracy,and Directional Accuracy Change are reported in percentage points (0.01 means 0.01%). For Rates, the ME is the difference $EstimatedRate-ActualRate$. MAE is calculated by taking the average of the absolute value of the ME. Directional Accuracy is the proportion of predictions that correctly identify whether the rate or level is above a threshold value (2.5% for GDP Growth, 3% for Inflation, 4% for Unemployment Rate, 4% for the 10-Year Treasury Yield, 16 for VIX, 1400 for Housing Starts, and 200 for Nonfarm Payrolls). For Levels, the MPE is calculated by taking the average of the percent error $(EstimatedLevel-ActualLevel)/ActualLevel$. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy Change is the proportion of predictions that correctly identify the direction of change (up or down) relative to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale from 0 to 100) and the MAPE. Num Obs is the number of observations used in the evaluation, Start Date and End Date indicate the period over which the metrics were computed. Refusals are the number of instances in which the model withheld a prediction by either answering ”null” or 0.
@@
Panel A: Rates ME (%) MAE (%) Directional Accuracy (%) Directional Accuracy Change (%) Confidence Calibration Start Date End Date Num Obs Refusals
Pre-cutoff
GDP Growth 0.01 0.15 96.27 96.99 -0.27 01/01/1990 07/01/2023 134 0
Inflation 0.00 0.04 98.02 93.07 -0.11 01/01/1990 09/01/2023 405 3
Unemployment Rate -0.00 0.03 99.26 83.46 0.09 12/01/1989 09/01/2023 406 0
10-Yr Treasury Yield -0.00 0.06 98.52 88.12 -0.40 01/31/1990 09/29/2023 405 0
Post-cutoff
GDP Growth -0.46 0.66 40.00 100.00 N/A 12/01/2023 12/01/2024 5 0
Inflation 0.35 0.38 47.06 56.25 0.70 10/01/2023 02/01/2025 17 0
Unemployment Rate -0.20 0.26 52.94 31.25 0.47 10/01/2023 02/01/2025 17 0
10-Yr Treasury Yield -0.26 0.49 27.78 47.06 0.14 10/31/2023 03/07/2025 18 0
@@
Panel B: Levels MPE (%) MAPE (%) Directional Accuracy (%) Directional Accuracy Change (%) Confidence Calibration Start Date End Date Num Obs Refusals
Pre-cutoff
VIX 10.42 13.62 100.00 83.94 -0.51 01/02/1990 09/29/2023 443 16
Housing Starts -2.38 3.93 100.00 81.86 -0.28 01/01/1990 09/01/2023 398 0
Nonfarm Payrolls -7.67 66.47 92.80 94.03 -0.11 01/01/1990 09/01/2023 403 1
Recent Pre-cutoff Period: Past 10 years
VIX 0.04 0.34 100.00 98.13 -0.20 10/31/2014 09/29/2023 108 0
Housing Starts -0.22 1.06 95.28 98.10 -0.14 10/01/2014 09/01/2023 106 0
Nonfarm Payrolls -0.00 0.00 100.00 100.00 -0.14 10/01/2014 09/01/2023 108 0
Post-cutoff
VIX 16.87 21.14 50.00 61.54 N/A 10/31/2023 02/28/2025 14 3
Housing Starts 2.25 8.54 35.29 56.25 N/A 10/01/2023 02/01/2025 17 0
Nonfarm Payrolls 69.51 97.44 58.82 68.75 -0.21 10/01/2023 02/01/2025 17 0
@@
This figure shows the LLM’s estimated values of some macro indicators compared to the actual values. Panels A, C, E, and G graph the actual values against the estimated values. Panels B, D, F, and H show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%).

### v2

@@ 3 Methodology
Our tests of LLMs’ recall capabilities span several categories of economic variables. First, we test macroeconomic indicators by querying precise values (e.g., unemployment rates) and directional trends. Second, we examine stock market indices through questions about exact numerical levels, directional changes, percentage movements, and relative performance. For example, we ask for the S&P 500 closing value on specific dates or whether the NASDAQ increased or decreased on particular days. Third, we test LLMs’ ability to identify news headlines’ dates by presenting sets of The Wall Street Journal front-page headlines without dates and asking it to identify when these headlines appeared and to predict the corresponding S&P 500 level on the next trading day. Fourth, we assess their ability to recall individual securities information, including specific stock price levels and directional movements. Fifth, we examine the effectiveness of requesting LLMs to impose artificial knowledge cutoff dates. Finally, we test the LLM’s ability to recover company identities from anonymized firm-specific text, including earnings calls and news headlines.
@@ 3.1 Period Selection and Supplementary Analysis
Our experimental design strategically spans three temporal zones relative to an LLM’s training cutoff date. First, we include periods before the cutoff where we expect high recall accuracy if memorization occurs. Second, we examine the periods within 10 years preceding the cutoff to assess potential recency effects in memorization patterns. Finally, we include post-cutoff periods as a control condition, where memorization is not expected.
@@ 3.2 Prompt
The context information represents the information set $x_{t}$ for our experimental conditions, which may be empty (when testing pure recall), contain historical data points, include relevant news headlines, or provide general knowledge about the period. More detailed information is provided in Section A of the Appendix.
For example, in a prompt testing recall with historical context, we might provide: “Context: The S&P 500 closed at 2,834.40 on March 14, 2019, and at 2,808.48 on March 13, 2019. What was the S&P 500 closing value on March 15, 2019?” This standardized approach allows us to systematically vary the information provided while controlling for confounding factors in question phrasing or response expectations.
3.3 LLMs
Given its wide usage in research, we conduct our main analysis using ChatGPT-4o with the specific version of “gpt-4o-2024-08-06.” This model is a snapshot of GPT-4o, ensuring consistent performance and behavior. It will not receive updates, and its training data ends in October 2023.777More information about this model is available here: https://platform.openai.com/docs/models/gpt-4o. Importantly, the temperature of the model is set to 0 for all of the analyses to maximize the reproducibility of the results. [^8] Temperature is a parameter of ChatGPT models that governs the randomness and the creativity of the responses. A temperature of 0 essentially means that the model will always select the highest probability word conditional on the text, which will eliminate the effect of randomness in the responses and maximize the reproducibility of the results. In additional analyses, we also test for the recall capabilities of open-source models such as Llama-3.1-70b-Instruct. [^9] Llama-3.1-70B-Instruct is Meta’s 70B-parameter instruction-tuned Llama-3.1 model, released July 23, 2024. More information is available here: https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct.
@@ 4 Data
Most of the tests use data from January 1990 to September 2023, which precedes GPT-4o’s knowledge cutoff date of October 2023. To test the LLM’s memorization, we use three categories of datasets: (1) stock index and individual stock prices, (2) macroeconomic indicators, and (3) textual data on WSJ front-page news, earnings calls, and firm-specific news headlines.
We ask the LLM to give us the closing value of the stock indices and a sample of individual stocks. We use the daily closing values of the S&P 500, the Dow Jones Industrial Average, and the Nasdaq Composite from Yahoo Finance to evaluate the LLM’s answers. We use Center for Research in Security Prices (CRSP) data to obtain daily stock market data for individual stock closing prices. The sample of individual stocks includes the Magnificent 7 (AAPL, AMZN, GOOGL, META, MSFT, NVDA, TSLA) and randomly selected subsamples of stocks in different size categories, representing a total of over 4,200 unique ticker symbols.
We also ask the LLM to give us estimates of various macroeconomic indicators. The indicators we test are (i) US GDP growth, (ii) inflation, (iii) unemployment rate, (iv) 10-year Treasury Yield, (v) VIX, (vi) housing starts, and (vii) change in nonfarm payrolls. We obtain the actual unemployment rate and 10-year Treasury Yield values from Federal Reserve Economic Data (FRED). We obtain the VIX levels from Yahoo Finance. For GDP growth, inflation, housing starts, and change in nonfarm payrolls, we use the Philadelphia Federal Reserve Real-Time Data Set to get the first vintage and ask the LLM to give us the earliest estimate of these indicators.
The textual data we use include The Wall Street Journal (WSJ) front-page headlines obtained from Factiva, earnings conference call transcripts from Capital IQ, firm-specific headlines, and the RavenPack news database. The WSJ front-page news dataset comprises 90,123 headlines from January 1990 to February 2025. There are, on average, approximately 9 headlines for each date. Given each set of headlines, we ask the LLM to provide the date and S&P 500 level on the next trading day. The conference call dataset starts in January 2006 and ends in May 2024. We extract the opening statement delivered by the CEO, anonymize the text using an entity neutering approach as proposed by [13], and ask the LLM to provide the firm, quarter, and year of the conference call.
We implement a similar test for the firm-specific headlines from January 2000 to June 2024. We follow the filtering procedures based on the RavenPack news database as in [22], including keeping only headlines with a relevance score of 100, keeping complete articles and press releases only and excluding headlines categorized as “stock-gain” and “stock-loss”, and avoiding repeated news through requiring the “event similarity days” to exceed 90 and removing duplicate headlines. We arrive at a sample of 1,358,737 headlines covering 4,391 unique firms from January 2000 to June 2024. About 70% of headlines are from overnight articles released either before 9 a.m. or after 4 p.m., with the remaining 30% from intraday articles.
@@ 5 Results
In this section, we present a comprehensive evaluation of GPT-4o’s memorization of economic and financial data, spanning macro indicators, market indices, individual stocks, headlines, and attempts to mitigate memorization through fake knowledge cutoffs and masking. Across these domains, we assess the model’s ability to recall precise values, identify contextual details, and adhere to constraints. Our goal is to study the extent and selectivity of memorization, highlighting its implications for using LLMs in economic forecasting and the challenges of isolating genuine predictive ability. Each subsection examines a specific data type or mitigation strategy, building a cohesive picture of how memorization manifests and persists.
First, we establish that memorization exists through direct elicitation tests on pre-cutoff data, where we query the model for specific values using only variable names and dates. High accuracy on these tests provides unambiguous evidence of memorization. Second, we examine post-cutoff performance as descriptive evidence regarding data contamination. When post-cutoff samples are small, hypothesis tests have low power to distinguish genuine forecasting skill from undetected memorization. Therefore, we interpret post-cutoff results cautiously: poor post-cutoff performance is descriptively consistent with absence of data contamination, but our post-cutoff samples are often small and do not constitute definitive statistical tests.
@@ 5.1 Macroeconomic Data Recall
To assess GPT-4o’s memorization of macroeconomic indicators, we tested its ability to recall monthly values across various variables (quarterly for GDP). In the baseline analysis, we use data from January 1990 to September 2023, all within the model’s training cutoff of October 2023. Importantly, for comparison, we also examine the post-cutoff period from October 2023 to February 2025. The indicators were divided into two groups: rates (GDP Growth, Inflation, Unemployment Rate, and 10-year Treasury Yield) and levels (Housing Starts, VIX, and Nonfarm Payrolls). For rates, we requested percentage values, evaluating accuracy through Mean Error, Mean Absolute Error, Threshold Accuracy (correctly identifying whether the rate was above a threshold value), and Directional Accuracy (correct direction of change from the previous period). For levels, we requested raw values, with errors calculated relative to actual levels through Mean Percent Error, Mean Absolute Percent Error, Threshold Accuracy, and Directional Accuracy. We also examined levels over the 10-year period preceding the cutoff to explore potential recency effects. We measure performance against the actual values of macro indicators and report the findings in Table 1.101010Table A1 in the Appendix reports summary statistics for the data to serve as benchmark accuracy rates for comparing the model’s ability to recall values against the true data.
@@
The results reveal a strong ability of GPT-4o to recall macroeconomic data. Figure 3 plots the recalled values against actual values as well as the estimation error for GDP Growth, Inflation, Unemployment Rate, and 10-year Treasury Yield. The series for recalled and actual values are almost identical, particularly in more recent periods. For rates, the model demonstrates near-perfect recall, with Mean Absolute Errors ranging from 0.03% (Unemployment Rate) to 0.15% (GDP Growth) and Threshold Accuracy exceeding 96% across all indicators, reaching 98% for 10-year Treasury Yield and 99% for Unemployment Rate. This set of results suggests that GPT-4o has memorized these percentage-based indicators with a high degree of accuracy.
@@
For levels, the recall remains high, with Threshold Accuracies between 94% and 100% for all indicators during the whole pre-training sample. Moreover, when focusing on the most recent 10-year period in the pre-training sample, performance improves dramatically–Mean Absolute Percent Errors fall to 1.06% for Housing Starts, 0.34% for VIX, and 0.00% for Nonfarm Payrolls, with Threshold Accuracy rising to 95%–100%. This recency effect indicates stronger memorization for more recent data, likely due to denser representation in the training corpus.
In contrast, the evaluation metrics in the post-cutoff period do not suggest recall. For rates, the Mean Absolute Errors increase significantly, ranging between 0.26% (Unemployment Rate) to 0.95% (GDP Growth). Threshold Accuracy falls to between 29.41% (10-Yr Treasury Yield) to 52.94% (Unemployment Rate). For levels, we see the same breakdown in recall. The Mean Absolute Percent Error jumps dramatically when compared to the recall in the more recent pre-cutoff period. For example, the Mean Absolute Percent Error rises from 0% to 97.4% for Nonfarm Payrolls. Threshold Accuracy drops from above 94% for all indicators to 35.3% (Housing Starts), 50.0% (VIX), or 58.8% (Nonfarm Payrolls).
@@
The high recall accuracy for rates and recent levels underscores the memorization problem when evaluating LLMs’ forecasting capabilities during the training period. The model’s ability to reproduce precise macroeconomic values, especially for percentage-based indicators and recent periods, suggests that apparent forecasting success for pre-cutoff data may stem from retrieving memorized information rather than genuine economic analysis. The weaker performance for levels over the full period, particularly for volatile indicators like Nonfarm Payrolls, hints at selective memorization, where certain data types or time frames are less reliably retained. On the other hand, we do not observe post-cutoff recall.
@@
This table reports a set of evaluation metrics for various macroeconomic indicators grouped into two panels: Rates and Levels. We ask the LLM to recall monthly values (quarterly for GDP, specific end of month date for 10-Year Treasury Yield and VIX) for each indicator. The indicators in the Rates panel include GDP Growth, Inflation, Unemployment Rate, and the 10-Year Treasury Yield. For these indicators, we ask the LLM to give us a percentage. The Levels panel includes Housing Starts, VIX, and Nonfarm Payrolls, evaluated over the full sample period. In each panel, we examine the periods before and after the knowledge cutoff date (October 2023) separately. The Levels, Recent Pre-cutoff Period: Past 10 years panel evaluates these same indicators over a more recent, shorter period. Mean Error (ME), Mean Absolute Error (MAE), Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), Threshold Accuracy,and Directional Accuracy are reported in percentage points (0.01 means 0.01%). For Rates, the ME is the difference EstimatedRateActualRateEstimatedRate-ActualRate. MAE is calculated by taking the average of the absolute value of the ME. Threshold Accuracy is the proportion of predictions that correctly identify whether the rate or level is above a threshold value (2.5% for GDP Growth, 3% for Inflation, 4% for Unemployment Rate, 4% for the 10-Year Treasury Yield, 16 for VIX, 1400 for Housing Starts, and 200 for Nonfarm Payrolls). For Levels, the MPE is calculated by taking the average of the percent error (EstimatedLevelActualLevel)/ActualLevel(EstimatedLevel-ActualLevel)/ActualLevel. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy is the proportion of predictions that correctly identify the direction of change (up or down) relative to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale from 0 to 100) and the MAPE. Num Obs is the number of observations used in the evaluation. Refusals are the number of instances in which the model withheld a prediction by either answering ”null” or 0. Refusal count also includes instances of missing data.
@@
Panel A: Rates ME (%) MAE (%) Threshold Accuracy (%) Directional Accuracy (%) Confidence Calibration Num Obs Refusals
Pre-cutoff, 01/1990 to 09/2023
GDP Growth 0.01 0.15 96.27 96.99 -0.27 134 1
Inflation 0.00 0.04 98.02 93.07 -0.11 405 0
Unemployment Rate -0.00 0.03 99.26 83.42 0.09 405 0
10-Yr Treasury Yield -0.00 0.06 98.52 88.12 -0.40 405 0
Post-cutoff, 10/2023 to 02/2025
GDP Growth 0.01 0.95 40.00 100.00 - 6 0
Inflation 0.35 0.38 47.06 56.25 0.70 17 0
Unemployment Rate -0.20 0.26 52.94 31.25 0.47 17 0
10-Yr Treasury Yield -0.24 0.48 29.41 50.00 0.16 17 0
@@
Panel B: Levels MPE (%) MAPE (%) Threshold Accuracy (%) Directional Accuracy (%) Confidence Calibration Num Obs Refusals
Pre-cutoff, 01/1990 to 09/2023
VIX 3.25 6.74 94.62 87.40 -0.42 390 15
Housing Starts -2.38 3.92 100.00 81.91 -0.24 399 6
Nonfarm Payrolls -7.65 66.30 95.56 94.06 -0.11 405 0
Recent Pre-cutoff Period, 10/2014 to 09/2023
VIX 0.04 0.34 100.00 98.13 -0.20 108 0
Housing Starts -0.22 1.06 95.28 98.10 -0.14 106 2
Nonfarm Payrolls -0.00 0.00 100.00 100.00 -0.14 108 0
Post-cutoff, 10/2023 to 02/2025
VIX 16.87 21.14 50.00 61.54 - 14 3
Housing Starts 2.25 8.54 35.29 56.25 - 17 0
Nonfarm Payrolls 69.51 97.44 58.82 68.75 -0.21 17 0
@@
This figure shows the LLM’s estimated values of macro indicators including Inflation, 10-yr Treasury Yield, GDP Growth, and Unemployment Rate compared to the actual values. Panels A, C, E, and G graph the actual values against the estimated values. Panels B, D, F, and H show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The post-cutoff period (10/2023 onward) is shaded gray.